# Refinement-based Christoffel sampling for least squares approximation in non-orthogonal bases

### v1

@@ 3.3 Discussion
For polynomial spaces, the growth of the Christoffel function on irregular domains has been studied extensively [27, 15]. Even if one does not have a tight upper bound, overestimating it has little impact since the computational complexity depends only logarithmically on this bound.
@@
A natural choice for an upper bound on the numerical dimension $n^{\epsilon}$ is $n$, since $n^{\epsilon}\leq n$ for all $\epsilon\geq 0$. In this case, the algorithm produces a sample set of size $\mathcal{O}(n\log(n))$, which suffices in most practical scenarios. If the basis $\Phi_{n}$ is highly numerically redundant, however, it can be advantageous to use a numerical dimension $n^{\epsilon}\ll n$.
@@ 4 Theoretical results
The algorithm presented in this paper is a continuous analogue of the iterative leverage score sampling algorithm proposed in [10, Algorithm 3], while incorporating $\ell^{2}$-regularization as in [11]. We heavily rely on this connection to prove the convergence behaviour of our algorithm.
The main goal is to prove that in each step of the algorithm, the updated function $u^{(new)}$ (7) remains an upper bound to $k_{n}^{\epsilon}$ while the $L^{1}_{\rho}(X)$-norm gets reduced compared to $u$. This is quantified in Theorem 4.1. For clarity, we are more explicit about the basis in our notation throughout this section. More specifically, we use
@@ 2. Establishing (10)
with probability at least $1-\gamma$ using a similar reasoning as in Lemma 3.1 for the basis $v\Phi$. More specifically, note that the samples used to construct $V$ are drawn i.i.d. from $\mu_{u}$. In order for (12) to hold, we need that the number of samples $\lceil\alpha m+1\rceil$ satisfies
@@
Due to lemma B.1, it suffices that
@@
$\leq\frac{m}{\lceil\alpha m+1\rceil}k^{\epsilon}\left(x;v\Phi,\frac{m}{\lceil\alpha m+1\rceil}V^{*}V\right)\qquad$ (Lemma B.3)
@@
where in the third step we use Lemma B.1. Since this bound holds for every $\delta_{2}>0$, one has
@@ 3. Establishing the probability
The proof of Theorem 4.1 relies heavily on the following lemma. This lemma asserts the existence of a weight function $v$ (with $0\leq v(x)\leq 1,\forall x\in X$) such that the numerical inverse Christoffel function $k_{n}^{\epsilon}$ associated with the weighted basis $v\Phi$ lies below a prescribed function $u$ (up to an arbitrarily small constant $\delta_{1}>0$). Moreover, the function $v$ does not distort the basis functions $\Phi$ too much, in the sense of (15). Note that without condition (15), condition (14) would be trivial to satisfy—for instance, by choosing $v=0$.
This lemma is closely related to discrete analogues, namely [10, Theorem 2] (for leverage scores) and [11, Lemma 6.2] (for ridge leverage scores with variable regularization parameter). On top of this connection, our proof proceeds by first discretizing the problem and then applying the proof techniques from [10, Theorem 2] and [11, Lemma 6.2]. Specifically, we require a discrete result on ridge leverage scores with a fixed regularization parameter, which is provided in Appendix A.
@@
Since $u$ is strictly positive, $u_{k}^{(\ell)}>0$ for all $i$. From Lemma A.1, it follows that there exists a diagonal matrix $V^{(\ell)}\in\mathbb{R}^{\ell\times\ell}$ with $0\preceq V^{(\ell)}\preceq I$ such that
@@ 1. Establishing (15)
for some $C>0$, where in the last step we use that the functions $\phi_{i}$ are Lipschitz continuous in every partition and bounded. Using [25, §21-1], it follows that there exists a permutation $\tau$ over $\{1,2,\dots,n\}$ such that
@@ 2. Establishing (14)
for some $C>0$, where in the last step we use that the functions $\phi_{i}$ are bounded and $\|M\|_{2}\leq\sqrt{n}\max_{1\leq i\leq n}\sum_{j=1}^{n}\lvert(M)_{ij}\rvert$ for $M\in\mathbb{C}^{n\times n}$ [25, §50-4]. This term also goes to zero as $h^{(\ell)}\to 0$ and, hence, the first error in (19) goes to zero as $h^{(\ell)}\to 0$.
@@
for some $C>0$, where in the last step we use that the function $u$ is Lipschitz continuous in $X_{k}$.
@@ 5 Numerical experiments
In this section we verify the performance of the refinement-based Christoffel sampling (RCS) algorithm numerically. Its implementation in Matlab and all code needed to reproduce the numerical experiments are made publicly available [22]. We use Matlab’s built-in function slicesample to sample from $\mu_{u}$, and we use either a simple Monte Carlo estimate or Matlab’s built-in function integral to estimate $\|u\|_{L_{\rho}^{1}(X)}$.
Some changes were made in the practical implementation compared to Algorithm 1, which favor the computational complexity while the empirical performance remains satisfactory. First, we discard $\log$-terms. More specifically, in each iterations we sample $C_{2}n^{\epsilon}$ points and construct the matrix $A$ via $(A)_{kj}=\overline{\phi_{j}(x_{k})}/\sqrt{C_{1}u(x_{k})}$. We choose $C_{1}=5$ and $C_{2}=5C_{1}$. At the end of the algorithm, we sample $C_{3}\|u\|_{L^{1}_{\rho}(X)}$ points with $C_{3}=10$. Second, we only take the first and the last two iterations into account for the update of $u$, i.e.,
@@
With these changes, the computational complexity reduces to $\mathcal{O}\left(n^{3}\log\left(\|k_{n}^{\epsilon}\|_{L^{\infty}(X)}\right)\right)$. For more details on the implementation, we refer to [22].
@@
    (21) $\Phi_{n}=\{p_{i}(x)\}_{i=0}^{n_{1}-1}\cup\{w(x)p_{i}(x)\}_{i=0}^{n_{2}-1},$
    where $n=n_{1}+n_{2}$ and $p_{i}$ is a polynomial basis function of degree $i$. Such bases are commonly used to approximate functions with known singular or oscillatory behaviour, which is incorporated in the weight function $w(x)$, see [4, Example 3] and references therein. The output of the RCS algorithm for $n=40$, X=[1,1]X=[-1,1] and $w(x)=\sqrt{x+1}$ is shown in Figure 2. In Figure 3, we consider the convergence of the least squares approximation of
    (22) $f(x)=\frac{\sqrt{x+1}}{1+5x^{2}}+\cos(5x)\qquad\text{for }x\in[-1,1]$
@@
    (a) linear scale
    (b) logarithmic scale
    Figure 2: Illustration of the RCS algorithm for the polynomial + weighted polynomial basis (21) with $w(x)=\sqrt{x+1}$ and $n_{1}=n_{2}=20$ ($n=40$). The function $u$ computed by the RCS algorithm is compared to an approximation of $k_{n}^{\epsilon}$ obtained via the dense grid method.
    Figure 3: Convergence of the discrete least squares approximation to $f$ (22) in the polynomial + weighted polynomial basis (21) with $w(x)=\sqrt{x+1}$. A comparison is made between approximation based on uniformly random samples and samples outputted by the RCS algorithm. The line indicates the geometric mean and the shaded area shows the variance over ten repetitions.
@@
    Rational approximation with preassigned poles. Recently, the lightning approximation scheme became popular for approximating functions with singularities at known locations, a setting which often occurs when solving PDEs on domains with corners [20]. As a model problem, consider approximation on $[0,1]$ of functions with an endpoint singularity at $x=0$ using the basis
    (23) $\Phi_{n}=\left\{\frac{-q_{i}}{x-q_{i}}\right\}_{i=1}^{n_{1}}\cup\{p_{i}(x)\}_{i=0}^{n_{2}-1}\qquad\text{with }q_{i}=-\exp\left(4\left(\sqrt{i}-\sqrt{n_{1}}\right)\right),$
    where $p_{i}$ is again a polynomial basis function of degree $i$, $n=n_{1}+n_{2}$ and $n_{2}=\mathcal{O}(\sqrt{n_{1}})$ [24]. Hence, we combine partial fractions with simple poles $q_{i}$, which cluster towards zero along the negative real axis in a tapered fashion [32], with a polynomial basis.
    Figure 5 shows the output of the RCS algorithm for different values of $n$. It becomes apparent that $k_{n}^{\epsilon}$ grows exponentially with the number of basis functions near the singularity. On Figure 5, the convergence behaviour of the discrete least squares fit to $f(x)=\sqrt{x}$ using $n_{2}=2\sqrt{n_{1}}$ is shown.
    Figure 4: Illustration of the RCS algorithm for the lightning + polynomial basis (23) for varying $n_{1}$ and $n_{2}$. The function $u$ computed by the RCS algorithm is compared to an approximation of $k_{n}^{\epsilon}$ obtained via the dense grid method.
    Figure 5: Convergence of the discrete least squares approximation to $f(x)=\sqrt{x}$ in the lightning + polynomial basis (23). A comparison is made between approximation based on uniformly random samples and samples outputted by the RCS algorithm. The line indicates the geometric mean and the shaded area shows the variance over ten repetitions.
    The lightning approximation scheme is extended to multiple dimensions in [8]. In order to illustrate the RCS algorithm in two dimensions, we consider the setting in [8, §4.2]. Namely,
@@
    Hence, we use $n=2n_{1}n_{2}^{2}+n_{3}^{2}$ basis functions. This basis is tailored to approximate functions with a curve of singularities along the elliptic curve $Q(x,y)=0$.
    In Figure 6, the output of the RCS algorithm is shown for $n_{1}=15$, $n_{2}=3$ and $n_{1}=10$ ($n=370$). The function $u$ is peaked along the elliptic curve, resulting in many clustered sample points close to the curve.
    (a) $u$
    (b) sample points
    Figure 6: Output of the RCS algorithm for the 2D lightning + polynomial basis (24) for $n_{1}=15,n_{2}=3$ and $n_{1}=10$ ($n=370$).
@@
    Fourier extension of a surface. Non-orthogonal bases often arise when approximating on irregular domains. In such cases, many approximation methods first construct an orthonormal basis on a regular domain. However, when this basis is restricted to the irregular domain, it generally loses orthogonality. In the case of a Fourier basis, this approach is referred to as Fourier extension [28].
    We consider the extreme case of using a 3D Fourier basis to approximate a function on the surface of a box. More specifically, we use
    (25) $\Phi_{n}=\{\exp(2\pi i(k_{x}x+k_{y}y+k_{z}z))\}_{(k_{x},k_{y},k_{z})=(-n_{x},-n_{y},-n_{z})}^{(n_{x},n_{y},n_{z})}$
    to approximate functions on the surface of a cube inside the torus $1/2,1/2)^{3}$. The output of the RCS algorithm is shown in Figure 7. We find that the function $u$ computed by the RCS algorithm peaks near the edges and corners of the cube.
    (a) $u$
    (b) sample points
    Figure 7: Output of the RCS algorithm for the Fourier basis (25) restricted to the surface of a cube for $n_{x}=n_{y}=n_{z}=4$ ($n=729$).

### v2

@@ 3.3 Discussion
For polynomial spaces, the growth of the Christoffel function on irregular domains has been studied extensively [kroo2015christoffel, dolbeault2022optimal]. Even if one does not have a tight upper bound, overestimating it has little impact since the computing time depends only logarithmically on this bound.
@@
A natural choice for an upper bound on the numerical dimension $n^{\epsilon}$ is $n$, since $n^{\epsilon}\leq n$ for all $\epsilon\geq 0$. In this case, the algorithm produces a sample set of size $\mathcal{O}(n\log(n))$, which suffices in most practical scenarios. If the set $\Phi_{n}$ is highly numerically redundant, however, it can be advantageous to use a numerical dimension $n^{\epsilon}\ll n$.
@@ 4 Theoretical results
In this section, we establish that the proposed algorithm converges with high probability. We begin by stating the main result—a corollary of Theorem 4.2, proved below, together with the algorithmic description given above.
Corollary 4.1.
Let $X\subseteq\mathbb{R}^{d}$ be compact, $\Phi=\{\phi_{i}\}_{i=1}^{n}\subset L^{2}_{\rho}(X)$ be a set of piecewise Lipschitz continuous functions, $\epsilon>0$ and assume $\inf_{x\in X}k^{\epsilon}(x)>0$. Run Algorithm 1 for some $0<\Delta<1$, upper bounds $K\geq\|k_{n}^{\epsilon}\|_{L^{\infty}}$ and $N\geq n^{\epsilon}$, sufficiently large $C_{1}$ and for some $C_{2}$ satisfying
$C_{2}>\frac{2C_{1}}{1-\Delta}\qquad\text{s.t.}\qquad C=\frac{(1-\Delta)C_{2}}{2C_{1}}>1.$
Then, with high probability (depending on $C_{1}$), the algorithm terminates in at most $\max\left(0,\left\lceil\log_{C}\left(\frac{(1-\Delta)K}{2N}\right)-1\right\rceil\right)$ iterations and produces at most $C_{2}N\log(N)$ samples and weights whose associated discrete Gram matrix satisfies (5), thereby ensuring accurate least squares fitting as described in corollaries 3.4 and 3.5.
The key step in proving this result is to show that, in each step of the algorithm, the updated function $u^{(new)}$ defined in (7) remains an upper bound for $k_{n}^{\epsilon}$, while its $L^{1}_{\rho}(X)$-norm decreases relative to that of $u$. This property is quantified precisely in Theorem 4.2. The algorithm presented in this paper can be viewed as a continuous analogue of the iterative leverage score sampling algorithm from [cohen2015uniform, Algorithm 3], while incorporating $\ell^{2}$-regularization as in [cohen2017input]. The analysis relies heavily on this connection.
For clarity, we use a more explicit notation throughout this section. More specifically, we use
@@ 2. Establishing (10)
with probability at least $1-\gamma$ using a similar reasoning as in Lemma 3.1 for the set $v\Phi$. More specifically, note that the samples used to construct $V$ are drawn i.i.d. from $\mu_{u}$. In order for (12) to hold, we need that the number of samples $\lceil\alpha m+1\rceil$ satisfies
@@
Due to Lemma SM2.1, it suffices that
@@
$\leq\frac{m}{\lceil\alpha m+1\rceil}k^{\epsilon}\left(x;v\Phi,\frac{m}{\lceil\alpha m+1\rceil}V^{*}V\right)\qquad$ (Lemma SM2.2)
@@
where in the third step we use Lemma SM1.1. Since this bound holds for every $\delta_{2}>0$, one has
@@ 3. Establishing the probability
The proof of Theorem 4.2 relies heavily on the following lemma, which asserts the existence of a weight function $v$ (with $0\leq v(x)\leq 1,\forall x\in X$) such that the numerical inverse Christoffel function $k_{n}^{\epsilon}$ associated with the weighted set $v\Phi$ lies below a prescribed function $u$ (up to an arbitrarily small constant $\delta_{1}>0$). Moreover, the function $v$ does not distort the functions in $\Phi$ too much, in the sense of (15). Note that without condition (15), condition (14) would be trivial to satisfy—for instance, by choosing $v=0$.
This lemma is closely related to discrete analogues, namely [cohen2015uniform, Theorem 2] (for leverage scores) and [cohen2017input, Lemma 6.2] (for ridge leverage scores with variable regularization parameter). On top of this connection, our proof proceeds by first discretizing the problem and then applying the proof techniques from [cohen2015uniform, Theorem 2] and [cohen2017input, Lemma 6.2]. Specifically, we require a discrete result on ridge leverage scores with a fixed regularization parameter, which is provided in the supplementary material.
@@
Since $u$ is strictly positive, $u_{k}^{(\ell)}>0$ for all $i$. From Lemma SM1.1, it follows that there exists a diagonal matrix $V^{(\ell)}\in\mathbb{R}^{\ell\times\ell}$ with $0\preceq V^{(\ell)}\preceq I$ such that
@@ 1. Establishing (15)
for some $C>0$, where in the last step we use that the functions $\phi_{i}$ are Lipschitz continuous in every partition and bounded. Using [hogben2006handbook, §21-1], it follows that there exists a permutation $\tau$ over $\{1,2,\dots,n\}$ such that
@@ 2. Establishing (14)
for some $C>0$, where in the last step we use that the functions $\phi_{i}$ are bounded and $\|M\|_{2}\leq\sqrt{n}\max_{1\leq i\leq n}\sum_{j=1}^{n}\lvert(M)_{ij}\rvert$ for $M\in\mathbb{C}^{n\times n}$ [hogben2006handbook, §50-4]. This term also goes to zero as $h^{(\ell)}\to 0$ and, hence, the first error in (19) goes to zero as $h^{(\ell)}\to 0$.
@@
for some $C>0$, where in the last step we use that $u$ is Lipschitz continuous in $X_{k}$.
@@ 5 Numerical experiments
In this section we verify the performance of the refinement-based Christoffel sampling (RCS) algorithm numerically. Its implementation in MATLAB and Julia and all code needed to reproduce the numerical experiments are made publicly available [github_repo]. We use slice sampling [neal2003slice] to (approximately) draw from $\mu_{u}$ and a simple Monte Carlo estimate to approximate $\|u\|_{L_{\rho}^{1}(X)}$.
Some changes were made in the practical implementation compared to Algorithm 1, which favor the computational complexity while the empirical performance remains satisfactory. First, we discard $\log$-terms. More specifically, in each iterations we sample $C_{2}n^{\epsilon}$ points and construct the matrix $A$ via $(A)_{kj}=\overline{\phi_{j}(x_{k})}/\sqrt{C_{1}u(x_{k})}$. We choose $C_{1}=5$ and $C_{2}=5C_{1}$. At the end of the algorithm, $C_{3}\|u\|_{L^{1}_{\rho}(X)}$ samples are drawn, with $C_{3}=10$ unless stated otherwise. By varying $C_{1}$, $C_{2}$ and $C_{3}$, one balances efficiency and reliability. Second, we only take the first and the last two iterations into account for the update of $u$, i.e.,
@@
With these changes, the computational complexity reduces to $\mathcal{O}\left(n^{3}\log\left(\|k_{n}^{\epsilon}\|_{L^{\infty}(X)}\right)\right)$. For more details on the implementation, we refer to the GitHub repository [github_repo].
We consider a couple of examples, which aim to showcase the versatility, robustness and efficiency of the algorithm. As a simple benchmark, we illustrate the improvement of RCS sampling compared to uniform sampling in Figure 3. Moreover, we compare the performance and efficiency of the RCS algorithm to the dense grid method [dolbeault2022optimal, adcock2020near, migliorati2021multivariate] in Figure 5.
@@
    $\Phi_{n}=\{p_{i}(x)\}_{i=0}^{n_{1}-1}\cup\{w(x)p_{i}(x)\}_{i=0}^{n_{2}-1},$
    where $n=n_{1}+n_{2}$ and $p_{i}$ is a polynomial basis function of degree $i$. Such bases are commonly used to approximate functions with known singular or oscillatory behaviour, which is incorporated in the weight function $w(x)$, see [adcock2019frames, Example 3] and references therein. The function $u$ computed by the RCS algorithm for $n=40$, X=[1,1]X=[-1,1] and $w(x)=\sqrt{x+1}$ is shown in Figure 2(a). In Figure 3(a), we consider the convergence of the discrete least squares approximation of
    $f(x)=\frac{\sqrt{x+1}}{1+5x^{2}}+\cos(5x)\qquad\text{on }X=[-1,1]$
@@
    Rational approximation with preassigned poles. The lightning scheme uses rational functions with preassigned poles to approximate functions with singularities at known locations, a setting which often occurs when solving PDEs on domains with corners [gopal2019solving]. As a simple model problem, we consider approximation on $[0,1]$ of functions with an endpoint singularity at $x=0$ using
    $\Phi_{n}=\left\{\frac{-q_{i}}{x-q_{i}}\right\}_{i=1}^{n_{1}}\cup\{p_{i}(x)\}_{i=0}^{n_{2}-1}\qquad\text{with }q_{i}=-\exp\left(4\left(\sqrt{i}-\sqrt{n_{1}}\right)\right),$
    where $p_{i}$ is again a polynomial basis function of degree $i$, $n=n_{1}+n_{2}$ and $n_{2}=\mathcal{O}(\sqrt{n_{1}})$ [herremans2023resolution]. Hence, we combine partial fractions with simple poles $q_{i}$, which cluster towards zero along the negative real axis in a tapered fashion [trefethen2021exponential], with a polynomial basis. Figure 2(b.1) shows the function $u$ computed by the RCS algorithm for different values of $n$. In Figure 3(b.1), we compare the error of discrete least squares approximation to $f(x)=\sqrt{x}$ using $n_{2}=2\sqrt{n_{1}}$ for RCS samples and uniformly random samples. Similarly, Figure 5 presents a comparison with the discrete grid method, showing that the RCS algorithm (using $C_{3}=15$) achieves higher accuracy with significantly less computation time.
    The lightning approximation scheme can be extended to multiple dimensions [boulle2024multivariate]. As a multivariate example, we consider the setting in [boulle2024multivariate, §4.2], namely:
@@
    Hence, the number of basis functions equals $n=2n_{1}n_{2}^{2}+n_{3}^{2}$. This basis is tailored to approximate functions with a curve of singularities along the elliptic curve $Q(x,y)=0$. In Figures 2(b.2) and 4(b.2), the output of the RCS algorithm is shown for $n_{1}=15$, $n_{2}=3$ and $n_{1}=10$ ($n=370$). The function $u$ is peaked along the elliptic curve. In Figure 3(b.2), we show the convergence behaviour of discrete least squares approximation to $f(x,y)=\lvert Q(x,y)\rvert=\lvert x^{3}-2x+1-y^{2}\rvert$.
@@
    Fourier extension of a curve. Non-orthogonal bases typically appear when approximating on irregular domains. Many methods, often implicitly, construct an orthonormal basis on a surrounding regular domain, which then loses orthogonality when restricted to the irregular one [adcock2020approximating]. For a Fourier basis, this is known as Fourier extension [matthysen2018function]. We consider the extreme case of approximating a function on a curve using a 2D Fourier basis. More specifically, we use
    (22) $\Phi_{n}=\{\exp(2\pi i(k_{x}x+k_{y}y))\}_{(k_{x},k_{y})=(-n_{1},-n_{1})}^{(n_{1},n_{1})}$
    to approximate functions on the edge of a triangle in the torus $1/2,1/2)^{2}$. The $n=(2n_{1}+1)^{2}$ functions are heavily redundant on the curve, meaning that the numerical dimension $n^{\epsilon}\ll n$. In our algorithm, we find that $N=10n_{1}+50=\mathcal{O}(\sqrt{n})$ works well as an upper bound for $n^{\epsilon}$. The function $u$ computed by the RCS algorithm is shown in Figure 2(c) for $n_{1}=15$. In Figure 3(c), we show the error of a discrete least squares fit to $f(x,y)=\sqrt{x^{2}+y^{2}}$. Note that Christoffel sampling does not seem to outperform uniform sampling, a phenomenon which has been analysed for 1D Fourier extensions [herremans2025sampling, adcock2023fast].